Sunday, August 25, 2013

Night Log 2013-08-24

SA: Paul
SO: Thea
Others: Johan, Simon, Keith


- a clear and photometric night with variable seeing which fluctuated between 1'' and 3''. had some technical issues at the beginning of the night, but all ran smoothly afterwards. A 4 degree temperature difference between the 2m and 30m sensors meant bad dome seeing and therefore mirror alignments.

- we obtained data for the following programs:
- 2013-1-RSA_UKSC_OTH-001 (P2)
- 2013-1-RSA_HET_UKSC_OTH-001 (P0) - BVIT
- 2013-1-IUCAA-004 (P3)
- 2013-1-RSA_DC_OTH-001 (P2)
- 2013-1-UW_OTH-001 (P2)
- 2013-1-UNC_RSA_RU-001 (P2)
- 2013-1-POL_RSA-001 (P3) (x3)

No comments:

Post a Comment